OpenStudy (anonymous):

M(–3, –1) is the midpoint of RS. If S has coordinates (4, 6), what are the coordinates of R?

OpenStudy (anonymous):

(–10, –8) (–10, 13) (11, 13) (11, –8)

OpenStudy (ledah):

use the midpoint formula: x = (x1 + x2) / 2 and y = (y1 + y2) / 2 just plug is the values for each set of coordinates example:x = -5 y = -1 x1 = 3 y1 = 3 -5 = (3 + x2) / 2 -10 = 3 + x2 -13 = x2 -1 = (3 + x2) / 2 -2 = 3 + x2 -5 = x2

OpenStudy (anonymous):

Ohh